### v1

# The State and Fate of Summarization Datasets

Abstract
Automatic summarization has consistently attracted attention, due to its versatility and wide application in various downstream tasks. Despite its popularity, we find that annotation efforts have largely been disjointed, and have lacked common terminology. Consequently, it is challenging to discover existing resources or identify coherent research directions. To address this, we survey a large body of work spanning 133 datasets in over 100 languages, creating a novel ontology covering sample properties, collection methods and distribution. With this ontology we make key observations, including the lack in accessible high-quality datasets for low-resource languages, and the field’s over-reliance on the news domain and on automatically collected distant supervision. Finally, we make available a web interface that allows users to interact and explore our ontology and dataset collection, [^1] https://github.com/edahanoam/Awesome-Summarization-Datasets as well as a template for summarization data card, which can be used to streamline future research into a more coherent body of work.
The State and Fate of Summarization Datasets
Noam Dahan  and Gabriel Stanovsky The Hebrew University of Jerusalem noam.dahan1@mail.huji.ac.il
1 Introduction
Summarization is the task of shortening a text while preserving the most important information it contains. This task has been explored for over 60 years Luhn (1958) yet it remains a longstanding challenge, as evidenced by the steadily increasing number of summarization datasets released every year (see Figure 1).
Despite the many research efforts in creating summarization datasets, we find they lack standardization and common terminology. This shortcoming makes it difficult to discover existing resources and bottlenecks, and to identify general directions for future research.
In this work, we conduct the most comprehensive survey of summarization datasets, covering a wide collection of 133 datasets across 104 languages. In Section 3, we develop a comprehensive ontology geared towards modern summarization datasets, positioning each dataset along seven different scales, including information about the samples (language, language modality, domain and summarization shape) as well as collection methods (annotation efforts and sources of supervision), and the manner in which the dataset is distributed.
Figure 1: There is a continuous increase in the number of summarization datasets being published. However, we find that the field is missing standardization and common terminology.
In Section 4 we show that our proposed ontology surfaces five major trends and gaps in the field. First, we find that the lack of common terminology thus far often leads to inconsistency and to the omission of important details when presenting datasets, making comparison challenging. An important example of this inconsistency is the traditional abstractive-extractive distinction, which we argue have gradually shifted away from a binary trait to a more nuanced spectrum.
Second, we observe that the field is over-reliant on the news domain, which acts as a double-edged sword – being both easy to obtain and prone to low quality. This is especially evident for low-resource languages, where the vast majority of datasets are based on news articles.
Figure 2: Our ontology for summarization datasets, accompanied by our annotations. The percentages in the language column indicate the proportion of datasets that support each language, where we count each multilingual dataset as multiple monolingual datasets. The arrows showcase the decision pathways for selecting specific datasets.
Third, we find that the datasets available for low-resource languages tend to be of subpar standard and especially challenging to discover. This is because many languages are currently supported only through multilingual datasets, which tend to compromise quality for diversity Urlana et al. (2022), while also making it harder to know which languages are annotated in which dataset.
Fourth, we find that most of the annotations were not specifically created to serve as summaries.
Fifth, we observe that copyright concerns, especially in the news domain have been changing data distribution, with many datasets move from publishing the data towards publishing scripts which reconstruct it from URLs. We argue that this practice is problematic, as URLs may change over time. Furthermore, a few seminal summarization datasets that were published publicly have been either taken down or are considered legally questionable due to copyright violations.
In Section 5 we publish two resources to standardize and help organize the vast research in the field. First, we make our dataset collection publicly available through a web UI which allows users to search datasets matching different criteria according to our terminology, for example, retrieving all manually created datasets in a certain language. This UI serves a central location where researchers can discover datasets, as well as identify new research directions. We welcome community contributions to keep this resource updated with newly published research.
Second, inspired by work on model cards Mitchell et al. (2019), we propose a standard summarization data card which covers the different items in our ontology as well as important properties. This can benefit future dataset developers in making their work more accessible, easy to discover and comparable against other datasets.
We hope that our work helps chart the way for future work in the field, as well enable researchers from various disciplines in exploring the wide range of existing summarization datasets.
To conclude we make the following contributions:
  • We formulate an ontology for summarization datasets, covering various axes regrading the samples, their collection and distribution.
  • We survey a wide range of 133 datasets covering datasets in 104 languages, showing that our ontology is rich enough to surface five important trends and challenges for future work, including the lack of terminology, the subpar quality of datasets in low-resource languages, and more.
  • We make available two resources – a web interface to interact with our survey and discover summarization datasets, and a formal summarization data card, which will help streamline future datasets.
2 Dataset Collection
Shape Example
Paragraph(s) Canadian authorities say she was being sought for extradition to the US, where the company is being investigated for possible violation of sanctions against Iran. Canada’s justice department said Meng was arrested in Vancouver on Dec. 1…
One Sentence John Higgins believes he has the ability and the form to win his fifth World Championship.
Span Earlier this year the art world made a rare discovery – a painting by Leonardo da Vinci. Only some 15 paintings by Leonardo still exist, including the “Mona Lisa”. So, this painting – known as “Salvator Mundi” or “Savior of the World” – is a truly remarkable find.
Highlights - Web site operator wants to publish encyclopedia about Harry Potter novels.
- Judge awards "Harry Potter" author J.K. Rowling and publisher \$6,750 in damages.
- Rowling says she has long planned to publish her own encyclopedia.
Table 1: The different shapes of summaries in our collection. Examples are taken from Cheng and Lapata (2016); Hermann et al. (2015), Lins et al. (2019), Narayan et al. (2018) and Fabbri et al. (2019).
Our survey covers 133 datasets collected over nearly 20 years, spanning from 2005 to 2024. The full list of references can be found in Appendix A.
Venues.
Most datasets are drawn from papers published in top-tier NLP, ML and AI venues and workshops. [^2] Venues: ACL, NAACL, EMNLP, LREC, EACL, SIGIR, AAAI, COLING, IEEE, NeurIPS, CoNLL, and TACL. In order to expand the supported languages, we also include datasets published through shared tasks and in language-specific journals and workshops.
Filtering.
We manually verify that each of the datasets is available for use, and exclude papers describing currently inaccessible resources. We include papers that explicitly focused on creating summarization datasets or that focused on summarization models and produced data as a by-product.
Collected Data.
For each dataset, we gather information on its properties, collection methods and distribution. This methodology allowed categorizing the current resources options for researchers, analyzing emerging trends and identifying the gaps in the field.
In order to analyze the current state of summarization datasets, we gathered additional information not covered in the ontology: the scale of data by number of samples, the levels of abstraction of the summaries and the data quality assessment methods applied to the collected data.
3 Ontology of Summarization Datasets
We propose a categorization of summarization datasets across seven orthogonal axes, and provide an ontology for each, depicted in Figure 2, and elaborated below. This ontology helps contextualize the extensive research in summarization and identifies missing types of datasets which could be addressed in follow up work.
3.1 Language and Language Modality
In the context of summarization, language choice entails the language used for input and output texts. We find a diverse set of datasets in 104 languages (see Appendix A).
In addition, we categorize each dataset into three categories of language modality:
Monolingual.
These datasets contain texts in a single language for both input and output.
Multilingual.
These datasets can be thought of as a few monolingual datasets together, as they contain document-summary pairs in various languages.
Cross-lingual.
In these datasets, the language of the output summary is different from the language of the input document.
3.2 Domain
Domain is the area of knowledge from which the documents are sourced. We identified nine domains that encompass the vast majority of datasets: News, which relies on articles from online platforms; Dialogue, which includes sources such as meeting transcripts, movie scenes, interviews, and customer service interactions; Scientific papers, most commonly paired with their abstract as summary; Legal documents such as government bills and patent documents; Encyclopedia entries, primarily based on Wikipedia; Opinions and arguments including reviews, editorials, and debates; Social media texts, which are typically more freeform; Literature such as books and poetry; and Instructional content, like "how-to" guides.
3.3 Summary Shape
We categorize the shape of the summaries in the datasets into four broad types, shown in Table 1.
Single-sentence summary.
These are summaries ranging from a few words (like titles) to a lengthy sentence, composed of several dozens of words. This shape is sometimes referred to as “extreme summarization” (Narayan et al., 2018) or “headline generation” (Napoles et al., 2012).
Highlights summary.
These summaries are constructed from a few, possibly independent sentences. For example, they could be bullet points (e.g., from news articles).
Paragraph summary.
These datasets contain output texts composed of one or more paragraphs of coherent free text. Examples of this shape include abstracts of scientific papers and sub-titles of news articles.
Span summary.
Built to accommodate extractive summarization, these datasets use spans of text directly taken from the document as summaries.
3.4 Sources for Supervision and Annotation Efforts
We categorize datasets by how much manual annotation effort they required. On opposite ends are datasets that were gathered in an Automatic way, mainly through scraping web resources, and human annotated datasets — in which annotators wrote the gold summaries. We also find semi-automatic datasets — in which human annotators are involved in the process but only later in the pipeline. Datasets created using this method typically provide annotators with summary options that they can correct, filter, or accept.
We argue that an additional often overlooked dimension is the objective for which the summaries were originally created, marked as Supervision in Figure 2.
Natural.
These datasets utilize summaries that were organically created. For example, highlights in news articles  Cheng and Lapata (2016); Hermann et al. (2015) and educational summaries published along text to help students Ladhak et al. (2020b).
Distant.
Similarly to the Natural category, these datasets use pre-existing texts as annotations. However, these texts were not originally created with the intention of serving as summaries, such as abstracts of scientific papers Cohan et al. (2018) or titles and subheadings of news articles Napoles et al. (2012).
Dedicated.
This category features datasets in which the summaries were written due to the creators’ efforts and did not exist beforehand. It includes mostly human annotated datasets, but also synthetically created ones.
3.5 Availability
The distribution of datasets depends on how accessible they are to the research community. We identify three main forms of data availability.
Publicly Available.
These datasets are published through a platform like Hugging Face.
URL-based reconstruction.
In this category, the data is not directly published; instead, a method to reconstruct it is provided. For example, a list of URLs to scrape as done in Grusky et al. (2018), or data preparation scripts as done in Kryściński et al. (2021).
Upon request.
This availability type, which is most restricting, usually requires filling a request form or contacting the authors in order to receive the data.
4 The State of Summarization Datasets
Figure 3: Every dot on the graph corresponds to a dataset in our collection. Each dot’s position indicates the percentage of unique uni-grams in the summaries that are not found in the source texts. Only a minority of the datasets report evaluation of the properties of the summaries.
We thoroughly examine 133 summarization datasets under the lens of our ontology. Figure 2 presents the quantitative dataset distribution as well as example categorization of several datasets. Following, we discuss five major observations regarding the current state and challenges in summarization datasets.
4.1 There is a Lack in Common Terminology
While summarization is a versatile and expressive task applicable in many scenarios, we find that summarization datasets are presented using inconsistent terminology and often lacking details. This makes it difficult for researchers to identify available resources and to effectively compare and contrast them.
For example, determining the shape of the summaries (such as paragraph or highlights) often required to access the dataset and examine samples; understanding how the summaries were created frequently involved visiting the original website from which the data was scraped; while datasets often do not mention the language in which they were annotated, in contrast with recent emphasis on explicit language disclosure Bender (2019).
All of these highlight the necessity of a shared ontology expressive enough to capture currently important aspects in summarization datasets, and guide dataset developers in reporting their work.
Abstractive vs. extractive summarization is no longer a useful distinction.
An important case where terminology is lacking is the distinction between extractive summarization, where models are required to select sentences from the source document and abstractive summarization, where generative models can also produce novel text.
With the advent of powerful generative LLMs, our survey indicates that these distinctions are no longer useful, with many of the surveyed datasets not specifying whether their annotations were either extractive or abstractive.
Instead, we argue that abstractiveness is a spectrum rather than a binary trait. To support this, in Figure 3 we plot the novel n-grams ratio for datasets in our collection. This metric, which we find to be the most commonly used for measuring levels of abstractions, measures the percentage of n-grams appearing in the summaries but not in the document. Formally Narayan et al. (2018):
# New n-grams# n-grams# New $n-grams\frac$$\text${\# New $\text{\# n-grams}$}divide start_ARG # New n-grams end_ARG start_ARG # n-grams end_ARG (1)
On this scale, 0% can be achieved by restricting the summary to contain only vocabulary from the source document and thus lean towards extractive. While at the other end of the scale, 40% of the 1-grams in the summaries were novel and thus leaning towards abstractive. Other less common metrics include Coverage and Density Grusky et al. (2018) as well as Abstractivity (ABS; Bommasani and Cardie, 2020). All of which are based on the overlap between the summary and the source document.
Figure 4: The rise in the number of languages supported by summarization resources was achieved mainly through multilingual datasets. Most languages in our survey are only supported through them.
4.2 News is the Most Popular Domain, Despite its Major Shortcomings
Roughly 50% of the datasets we surveyed use news articles as source documents. This popularity may be due to the large number of online news platforms in various languages that allow for easy scraping. Moreover, many of these articles already include human-written summaries or distant supervision, in the form of highlights, titles, and subtitles. We argue that this wide availability and relative ease of obtaining distant supervision is in fact a double-edged sword, harming the quality of the gold summaries, while presenting a rather simple challenge for summarization models.
First, the news domain may present a simple instance of the summarization task. News articles typically follow “inverted pyramid writing style” — i.e., the most important information answering the 5 w’s (who? what? etc.) appears in the first few sentences (Pottker, 2003). Consequently, a summarization model which trivially outputs the first sentences of the news article can often achieve good results (Cohen et al., 2021).
Second, news summarization datasets are prone to be of low quality, because they often rely on automatic heuristics. Tejaswin et al. (2021) manually evaluated two datasets with one sentence summaries, one which uses the title as summary Rush et al. (2015), and another which uses a bold sentence in the beginning of the article Narayan et al. (2018). They found that about half of the samples are either incomplete or contain concepts or entities that do not appear in the rest of the document, and hence often cannot be deduced from it.
Beyond the news domain, the scientific domain, characterized by longer texts and complex terminology is popular Cohan et al. (2018), yet it remains predominantly English-centric. Official documents, such as those in the legal domain, patents, and government bills (e.g., Sharma et al., 2019; Kornilova and Eidelman, 2019; Huang et al., 2021), share similar complexity in that salient information is usually distributed throughout the document, and are more multilingual. We also find that dialogues are increasingly attracting the attention of the research community, with this domain being the second most popular, and that Wikipedia articles are a valuable resource due to their guidelines and API, as well as for cross-lingual research because they cover the same topics in multiple languages.
4.3 Datasets for Low-Resource Languages are Often Low Quality and Hard to Discover
The attention to performing NLP research on a diverse set of languages has been increasing in recent years (Joshi et al., 2020; Ruder, 2022). Evidently, while English is by far the most popular language in our survey, we cover resources for 104 different languages (see Appendix A for the complete list). We find that this encouraging versatility was achieved mainly through multi-lingual datasets, as 81 of the languages surveyed are supported solely through them. In Figure 4, we illustrate the growth in the number of languages with available resources over the years, emphasizing the dramatic rise following the introduction of multilingual resources.
However, while providing important diversity, we identify limitations in how multilingual summarization datasets are created and reported.
Discovering existing datasets in low-resource languages proves challenging.
Given that the majority of languages in our survey are available only within large multilingual corpora,
researchers must comb through all multilingual datasets in search for their language. This is challenging as there is currently no unified platform that houses all available datasets and allows to search by the desired language.
Figure 5: The distribution of domains for English datasets versus non-English datasets. While English domains are diverse, our survey shows most datasets for other languages are comprised of news articles.
To make matters worse, determining which languages are present in each multilingual dataset proved challenging, as we mentioned before, datasets often do not explicitly list the supported languages in the paper.
Datasets in low-resource languages are less diverse in terms of domain, and of poorer quality.
We find that the over-reliance on the news domain is considerably worse in non-English datasets. In Figure 5, we compare the domain distribution between English and non-English resources. We find that more than 70% of datasets for these languages are from the news domain. This exacerbates the problems we observed regarding the news domain. Indeed, Urlana et al. (2022) reported that 97% of the Telugu samples in XL-SUM Hasan et al. (2021) and MassiveSum Varab and Schluter (2021) were of low quality, with similar results in Hindi, Gujarati and Marathi.
4.4 Most Annotations were not Intended to be Used as Summaries
We found only 35 dedicated datasets, i.e., where the annotations were created specifically for the task and did not exist before. This is three times fewer than datasets that include existing annotations.
Not only that the vast majority of annotations available were gathered automatically, but datasets collected this way are an order of magnitude bigger: human annotated datasets were on average more than 10 times smaller in number of samples than automatically collected datasets and thus allow for training big models.
We categorized 43% of the datasets as Distant supervision, meaning the human authors did not explicitly intend to write a summary of the source document. This includes proxies of summaries like titles and subtitles in news articles, abstracts of scientific papers or descriptions of videos or interviews. Other approaches we consider as distant are social media metadata as summaries, as introduced in Newsroom Grusky et al. (2018) and beginning with summaries and searching for longer texts to support them Ghalandari et al. (2020); Saxena and Keller (2024).
We also find a mismatch between the datasets we create and the ones we use. In Figure 6, we show a clear trend over the years of collecting the paragraph shape. This might be considered a natural shape for summarization as it is also the most common for collecting data through human annotation in our collection. This comes in contrast to Highlights being very commonly used in research. This is evident from CNN-Daily Mail, perhaps the most popular summarization dataset at the time of writing, which uses this shape.
Figure 6: There is a clear trend of collecting summaries of paragraph shape. We note some datasets offer a few shapes.
4.5 Copyright Concerns Limit the Distribution of Datasets
Our survey shows that there is a recent trend away from publishing datasets in a public manner, instead making them available through reconstruction scripts from URLs (Grusky et al., 2018; Varab and Schluter, 2020; Kryściński et al., 2021), which may be due to growing legal concerns (Karamolegkou et al., 2023).
Reconstruction may be a temporary solution, which does not address the underlying legal issues. Moreover, this practice may lead to inconsistencies between local instances of the same dataset obtained at different times, due to changes in URL contents.
Recently, even datasets which were previously publicly available are being called into question. For example, the popular CNN/Daily Mail and XSUM are considered to be “legally questionable” Wang et al. (2022a). In a similar vein, the New York Times Corpus Sandhaus (2008), despite being published by the newspaper itself, was recently withdrawn and is no longer available. This may be another downside of the field’s over-reliance on the news domain.
5 Standardizing Summarization Datasets
Summarization Data Card
Sample information:
Languages:
List all supported languages
Summary Shape:
Paragraph/One Sentence/Highlights/Span
Domain:
Example: News/Scientific/Dialogues/etc.
Size:
Number of document-summary pairs
Annotation information:
Annotation efforts: Automatic, Human annotations, Semi-automatic
Source of supervision: Natural (summaries created organically)/ Distant (annotations are proxies of summaries)/ Dedicated (annotations created by researchers)
Brief description of the summaries’ source: Example: digests of legal documents
Data quality assessment:
Abstraction level: 1-to-4-gram ratios
[…]
Human evaluation: Yes/No
Availability details:
How is the data made accessible: Publicly Available / URL-based Reconstruction / Upon Request
Copyrights information: License
Table 2: Template for a summarization data card.
Our survey surfaces a lack in terminology and reporting of the vast number of summarization datasets. This challenges both users of datasets, who may find it hard knowing whether there exists a dataset specifically tailored to their needs (e.g., in terms of language, domain, or shape), as well as researchers in the field, looking to identify current bottlenecks and directions for future research. Here we outline two contributions we make towards making datasets more standard and accessible.
Interactive web interface.
We will make available a web UI populated with our ontology of 133 summarization datasets, along with our annotations of the seven axes which characterize each dataset (See Figure 7). Users of this interface can search for a dataset by specifying the criteria they are looking for. We will welcome community contributions to add new datasets, or past datasets not already included in it. We hope that this interface can become a central hub, allowing users to easily find relevant datasets, without the need to manually sift through hundreds of papers.
Template for a summarization data card.
In recent years, model and data cards have helped set agreed reporting standards, formalizing what details are reported when publishing a new resource (Mitchell et al., 2019). Moving forward, we suggest that summarization research adopt a similar approach. In Table 2 we propose a template for a summarization data card which follows findings from our survey.
Specifically, we focus on four key elements to report: (1) sample information: shape, languages, domain; (2) annotation information, indicating how were the summaries created: e.g., via scraping or manual effort; (3) data quality assessment: we propose to report the level of abstraction via n-gram novelty, the Compression rate (Grusky et al., 2018), as well as human evaluation of the data, if it was performed; and (4) availability: indicating copyright licensing and download information.
6 Related Work
Several surveys have explored different aspects of automatic summarization Kryściński et al. (2019); Cajueiro et al. (2023); Sharma and Sharma (2022), yet their focus was more on models rather than datasets. To the best of our knowledge, none of the previous work has focused on the state of low-resource languages, nor looked at the dataset collection method and distribution. Dernoncourt et al. (2018) published a repository of available corpora for summarization. However, it included only a limited number of datasets (21), with just four that are not in English.
Several surveys focus on a sub-task of summarization: Koh et al. (2022) reviewed long document summarization; Jangra et al. (2023) studied multimodal summarization; and Tuggener et al. (2021) conducted a survey of dialogue summarization datasets.
A few papers focused on assessing the properties of popular datasets. Bommasani and Cardie (2020) conducted evaluation of 10 famous summarization datasets using novel metrics and Tejaswin et al. (2021) manually analyse 600 samples from three datasets.
7 Conclusion
We carried out an extensive survey of 133 summarization datasets in 104 languages, introducing a new ontology covering sample information, collection methods and data distribution. We offer a detailed report on the current state of resources in the field, surfacing challenges due to the lack of terminology, as well as findings on low resource languages, data quality and availability. To support researchers, we release two resources: a platform to search available datasets by our categorization and a template for a summarization data card, to help streamline future datasets.
8 Limitations
Although this survey provides a thorough overview of the available summarization datasets, we wish to acknowledge several limitations. First, despite our best efforts, we may not have found all available resources. This is especially true for low resource languages and datasets published by less-established conferences.
Second, we decided to omit from our ontology some axis. Specifically, we did not analysed data about sub-tasks of summarizations such as multi-modal, long-form or aspect based. We note that the available resources are still included in our platform and marked accordingly.
Acknowledgments
We would like to thank Omer Kidron, Daria Lioubashevski and the members of the SLAB group at the Hebrew University for the productive discussion.
Appendix A Datasets Categorized by Language
In this section we list all datasets included in this survey per language.
Afrikaans:
Varab and Schluter (2021), Giannakopoulos et al. (2017)
Albanian:
Varab and Schluter (2021), Palen-Michel and Lignos
Amharic:
Hasan et al. (2021), Palen-Michel and Lignos , Bhattacharjee et al. (2021), Shi (2023), Varab and Schluter (2021), Liang et al. (2022), Bhattacharjee et al. (2021), Xiaorui (2023)
Arabic:
Giannakopoulos et al. (2015), El-Haj et al. (2010), Hasan et al. (2021), Nguyen and Daumé III (2019), El-Haj and Koulali (2013), Alhamadani et al. (2022), Chouigui et al. (2021), Varab and Schluter (2021), Liang et al. (2022), Giannakopoulos et al. (2017), Ladhak et al. (2020a), Giannakopoulos et al. (2017), Bhattacharjee et al. (2021), Tikhonov and Malykh (2022), Xiaorui (2023)
Armenian:
Palen-Michel and Lignos , Varab and Schluter (2021)
Assamese:
Urlana et al. (2023), Goutom et al. (2023), Varab and Schluter (2021)
Aymara:
Varab and Schluter (2021)
Azerbaijani:
Hasan et al. (2021), Palen-Michel and Lignos , Varab and Schluter (2021), Liang et al. (2022), Giannakopoulos et al. (2017),Bhattacharjee et al. (2021), Xiaorui (2023)
Bambara:
Varab and Schluter (2021)
Basque:
Giannakopoulos et al. (2017)
Bengali:
Hasan et al. (2021), Urlana et al. (2023), Palen-Michel and Lignos , Nguyen and Daumé III (2019), Khan et al. (2023), Varab and Schluter (2021), Liang et al. (2022), Verma et al. (2023), Bhattacharjee et al. (2021), Xiaorui (2023)
Bislama:
Varab and Schluter (2021)
Bosnian:
Palen-Michel and Lignos , Varab and Schluter (2021), Cao et al. (2020), Giannakopoulos et al. (2017)
Bulgarian:
Varab and Schluter (2021), Giannakopoulos et al. (2017), Aumiller et al. (2022)
Burmese:
Hasan et al. (2021), Palen-Michel and Lignos , Varab and Schluter (2021), Liang et al. (2022), Bhattacharjee et al. (2021), Xiaorui (2023)
Catalan:
Soriano et al. (2022), Varab and Schluter (2021), Giannakopoulos et al. (2017)
Central Khmer:
Palen-Michel and Lignos , Varab and Schluter (2021)
Chinese:
 [^3] This list includes both Traditional Chinese and Simplified Chinese
Giannakopoulos et al. (2015), Hu et al. (2015), Hasan et al. (2021), Palen-Michel and Lignos , Wu et al. (2023), Liu et al. (2020), Liu et al. (2022), Zang et al. (2022), Chen et al. (2021a), Zhao et al. (2021), Varab and Schluter (2021), Li et al. (2017), Liang et al. (2022), Lin et al. (2021), Verma et al. (2023), Li et al. (2020), Wang et al. (2021), Giannakopoulos et al. (2017), Ladhak et al. (2020a), Zhu et al. (2019), Wang et al. (2022b), Bhattacharjee et al. (2021), Takeshita et al. (2024a), Tikhonov and Malykh (2022), Xiaorui (2023)
Croatian:
Varab and Schluter (2021), Cao et al. (2020), Giannakopoulos et al. (2017), Aumiller et al. (2022)
Czech:
Giannakopoulos et al. (2015),Straka et al. (2018), Varab and Schluter (2021), Giannakopoulos et al. (2017), Ladhak et al. (2020a), Perez-Beltrachini and Lapata (2022), Aumiller et al. (2022)
Danish:
Varab and Schluter (2020), Varab and Schluter (2021), Aumiller et al. (2022)
Dari:
Palen-Michel and Lignos , Varab and Schluter (2021)
Dutch:
Nguyen and Daumé III (2019), Varab and Schluter (2021), Giannakopoulos et al. (2017), Ladhak et al. (2020a), Aumiller et al. (2022), Tikhonov and Malykh (2022)
English:
Giannakopoulos et al. (2015), Cheng and Lapata (2016); Hermann et al. (2015), Grusky et al. (2018), Hasan et al. (2021), Fabbri et al. (2019), Cachola et al. (2020), Takeshita et al. (2024b), Jiang and Dreyer (2024), Saxena and Keller (2024), Mahbub et al. (2023), Urlana et al. (2023), Datta et al. (2023), Klaus et al. (2022), Palen-Michel and Lignos , Cohen et al. (2021), Koupaee and Wang (2018), Huang et al. (2021), Sharma et al. (2019), Napoles et al. (2012), Nguyen and Daumé III (2019), Kornilova and Eidelman (2019), Hasler et al. (2003), Lins et al. (2019), Zhang and Tetreault (2019), Wang and Ling (2016), Ganesan et al. (2010), Cohan et al. (2018), Kim et al. (2018), Gliwa et al. (2019), Zhu et al. (2021), Yasunaga et al. (2019), Cao et al. (2016), Wei and Gao (2018), Narayan et al. (2018), Abacha and Demner-Fushman (2019), Yadav et al. (2022), Xu et al. (2021), Zhang and Wan (2017), Ghalandari et al. (2020), Sotudeh et al. (2021), Hayashi et al. (2021), Malykh et al. (2020), Ta et al. (2023), Syed et al. (2020), Liu et al. (2023), Kraaij et al. (2005), Völske et al. (2017), Chen et al. (2021c), Gupta et al. (2021), Kryściński et al. (2021), Ladhak et al. (2020b), Hayashi et al. (2020), Antognini and Faltings (2020), Wang et al. (2022a), Feigenblat et al. (2021), Lu et al. (2020), Chen et al. (2021b), Meng et al. (2021), Khalman et al. (2021), Roush and Balaji (2020), Zhong et al. (2021), Bhatia et al. (2014), Ulrich et al. (2008), Varab and Schluter (2021), Li et al. (2018), Li et al. (2017), Liang et al. (2022), Zhu et al. (2018), Chen and Zhuge (2018), Yadav et al. (2021), Gorinski and Lapata (2015), Verma et al. (2023), Sanabria et al. (2018), Wang et al. (2021), Zopf (2018), Zopf et al. (2016), Collins et al. (2017), Lev et al. (2019), Park and Park (2023), Giannakopoulos et al. (2017), Ladhak et al. (2020a), Zhu et al. (2019), Perez-Beltrachini and Lapata (2022), Wang et al. (2022b), Aumiller et al. (2022), Bhattacharjee et al. (2021), Takeshita et al. (2024a), Fatima and Strube (2021), Tikhonov and Malykh (2022), Xiaorui (2023)
Esperanto:
Varab and Schluter (2021), Giannakopoulos et al. (2017)
Estonian:
Aumiller et al. (2022)
Filipino:
Varab and Schluter (2021)
Finnish:
Giannakopoulos et al. (2017), Aumiller et al. (2022)
French:
Giannakopoulos et al. (2015), Hasan et al. (2021), Palen-Michel and Lignos , Scialom et al. (2020), Nguyen and Daumé III (2019), Varab and Schluter (2021), Liang et al. (2022), Eddine et al. (2020), Verma et al. (2023), Wang et al. (2021), Giannakopoulos et al. (2017), Ladhak et al. (2020a), Perez-Beltrachini and Lapata (2022), Bhattacharjee et al. (2021), Tikhonov and Malykh (2022), Xiaorui (2023)
Fulah:
Varab and Schluter (2021)
Georgian:
Palen-Michel and Lignos , Varab and Schluter (2021), Giannakopoulos et al. (2017)
German:
Scialom et al. (2020), Nguyen and Daumé III (2019), Varab and Schluter (2021), Kew et al. (2023), Wang et al. (2021), Zopf (2018), Giannakopoulos et al. (2017), Wang et al. (2022b), Aumiller et al. (2022), Takeshita et al. (2024a), Fatima and Strube (2021), Tikhonov and Malykh (2022)
Gujarati:
Hasan et al. (2021), Satapara et al. (2022), Urlana et al. (2023), Varab and Schluter (2021), Liang et al. (2022), Verma et al. (2023), Ladhak et al. (2020a), Perez-Beltrachini and Lapata (2022), Bhattacharjee et al. (2021), Xiaorui (2023)
Haitian:
Palen-Michel and Lignos , Varab and Schluter (2021)
Hausa:
Hasan et al. (2021), Palen-Michel and Lignos , Varab and Schluter (2021), Liang et al. (2022), Bhattacharjee et al. (2021), Xiaorui (2023)
Hebrew:
Giannakopoulos et al. (2015), Varab and Schluter (2021), Mondshine et al. (2024)
Hindi:
Giannakopoulos et al. (2015), Hasan et al. (2021), Satapara et al. (2022), Urlana et al. (2023), Datta et al. (2023), Varab and Schluter (2021), Liang et al. (2022), Verma et al. (2023), Wang et al. (2021), Ladhak et al. (2020a), Bhattacharjee et al. (2021), Xiaorui (2023)
Hungarian:
Varab and Schluter (2021), Barta et al. (2024), Aumiller et al. (2022)
Icelandic:
Varab and Schluter (2021)
Igbo:
Hasan et al. (2021), Varab and Schluter (2021), Liang et al. (2022), Bhattacharjee et al. (2021), Xiaorui (2023)
Indonesian:
Hasan et al. (2021), Palen-Michel and Lignos , Koto et al. (2020), Kurniawan and Louvan (2018), Varab and Schluter (2021), Liang et al. (2022), Verma et al. (2023), Wang et al. (2021), Giannakopoulos et al. (2017), Ladhak et al. (2020a), Bhattacharjee et al. (2021), Xiaorui (2023)
Irish:
Varab and Schluter (2021), Aumiller et al. (2022)
Italian:
Landro et al. (2022), Nguyen and Daumé III (2019), Varab and Schluter (2021), Giannakopoulos et al. (2017), Ladhak et al. (2020a), Aumiller et al. (2022), Takeshita et al. (2024a), Tikhonov and Malykh (2022)
Japanese:
Hasan et al. (2021),Nguyen and Daumé III (2019), Varab and Schluter (2021), Liang et al. (2022), Yamamura and Shimada (2018), Verma et al. (2023), Giannakopoulos et al. (2017), Ladhak et al. (2020a), Bhattacharjee et al. (2021), Takeshita et al. (2024a), Tikhonov and Malykh (2022), Xiaorui (2023)
Javanese:
Giannakopoulos et al. (2017)
Kannada:
Urlana et al. (2023), Varab and Schluter (2021)
Kinyarwanda:
Palen-Michel and Lignos , Varab and Schluter (2021)
Kirghiz:
Hasan et al. (2021), Varab and Schluter (2021), Liang et al. (2022), Bhattacharjee et al. (2021), Xiaorui (2023)
Korean:
Hasan et al. (2021), Palen-Michel and Lignos , Liang et al. (2022), Giannakopoulos et al. (2017), Ladhak et al. (2020a), Bhattacharjee et al. (2021), Xiaorui (2023)
Kurdish:
Palen-Michel and Lignos , Badawi (2023), Varab and Schluter (2021)
Lao:
Palen-Michel and Lignos , Varab and Schluter (2021)
Latvian:
Varab and Schluter (2021), Giannakopoulos et al. (2017), Aumiller et al. (2022)
Limburgish:
Giannakopoulos et al. (2017)
Lingala:
Varab and Schluter (2021)
Lithuanian:
Varab and Schluter (2021), Aumiller et al. (2022)
Macedonian:
Palen-Michel and Lignos , Nguyen and Daumé III (2019), Varab and Schluter (2021)
Malagasy:
Nguyen and Daumé III (2019), Varab and Schluter (2021)
Malay:
Urlana et al. (2023), Giannakopoulos et al. (2017)
Malayalam:
Urlana et al. (2023), Varab and Schluter (2021)
Maltese:
Aumiller et al. (2022)
Manipuri:
Urlana et al. (2023)
Marathi:
Hasan et al. (2021), Urlana et al. (2023), Urlana et al. (2023), Varab and Schluter (2021), Liang et al. (2022), Verma et al. (2023), Giannakopoulos et al. (2017), Bhattacharjee et al. (2021), Xiaorui (2023)
Modern Greek:
Giannakopoulos et al. (2015), Palen-Michel and Lignos , Nguyen and Daumé III (2019), Varab and Schluter (2021), Giannakopoulos et al. (2017), Aumiller et al. (2022)
Mongolian:
Varab and Schluter (2021)
Nepali:
Hasan et al. (2021), Varab and Schluter (2021), Liang et al. (2022), Verma et al. (2023), Bhattacharjee et al. (2021), Xiaorui (2023)
North Ndebele:
Palen-Michel and Lignos , Varab and Schluter (2021)
Norwegian:
Giannakopoulos et al. (2017)
Oriya:
Urlana et al. (2023), Varab and Schluter (2021)
Oromo:
Hasan et al. (2021), Varab and Schluter (2021), Liang et al. (2022), Bhattacharjee et al. (2021), Xiaorui (2023)
Panjabi:
Hasan et al. (2021), Urlana et al. (2023), Varab and Schluter (2021), Liang et al. (2022), Verma et al. (2023), Bhattacharjee et al. (2021), Xiaorui (2023)
Pashto:
Hasan et al. (2021), Palen-Michel and Lignos , Varab and Schluter (2021), Liang et al. (2022), Verma et al. (2023), Bhattacharjee et al. (2021), Xiaorui (2023)
Persian:
Hasan et al. (2021), Palen-Michel and Lignos , Liang et al. (2022), Farahani et al. (2021), Giannakopoulos et al. (2017), Bhattacharjee et al. (2021), Xiaorui (2023)
Polish:
Varab and Schluter (2021), Giannakopoulos et al. (2017), Aumiller et al. (2022), Tikhonov and Malykh (2022)
Portuguese:
Hasan et al. (2021), Palen-Michel and Lignos , Nguyen and Daumé III (2019), Cardoso et al. (2011), Paiola et al. (2024), Varab and Schluter (2021), Liang et al. (2022), Verma et al. (2023), Sanabria et al. (2018), Wang et al. (2021), Giannakopoulos et al. (2017), Ladhak et al. (2020a), Aumiller et al. (2022), Bhattacharjee et al. (2021), Tikhonov and Malykh (2022), Xiaorui (2023)
Romanian:
Giannakopoulos et al. (2015), Varab and Schluter (2021), Giannakopoulos et al. (2017), Aumiller et al. (2022)
Rundi:
Hasan et al. (2021), Varab and Schluter (2021), Liang et al. (2022), Bhattacharjee et al. (2021), Xiaorui (2023)
Russian:
Hasan et al. (2021), Palen-Michel and Lignos , Scialom et al. (2020), Nguyen and Daumé III (2019), Gusev (2020), Varab and Schluter (2021), Liang et al. (2022), Verma et al. (2023), Wang et al. (2021), Giannakopoulos et al. (2017), Ladhak et al. (2020a), Bhattacharjee et al. (2021), Tikhonov and Malykh (2022), Xiaorui (2023)
Scottish Gaelic:
Hasan et al. (2021), Varab and Schluter (2021), Liang et al. (2022), Bhattacharjee et al. (2021), Xiaorui (2023)
Serbian:
 [^4] This list includes Cyrillic and Latin writing systems
Hasan et al. (2021), Palen-Michel and Lignos , Varab and Schluter (2021), Liang et al. (2022), Bhattacharjee et al. (2021), Xiaorui (2023)
Shona:
Palen-Michel and Lignos , Varab and Schluter (2021)
Sinhala:
Hasan et al. (2021), Varab and Schluter (2021), Liang et al. (2022), Verma et al. (2023), Bhattacharjee et al. (2021), Xiaorui (2023)
Slovak:
Ondrejova and Šuppa (2024), Šuppa and Adamec (2020), Varab and Schluter (2021), Giannakopoulos et al. (2017), Aumiller et al. (2022)
Slovenian:
Varab and Schluter (2021), Aumiller et al. (2022)
Somali:
Hasan et al. (2021), Palen-Michel and Lignos , Varab and Schluter (2021), Liang et al. (2022), Bhattacharjee et al. (2021), Xiaorui (2023)
Spanish:
Giannakopoulos et al. (2015), Hasan et al. (2021), Soriano et al. (2022), Palen-Michel and Lignos , Scialom et al. (2020), Nguyen and Daumé III (2019), Varab and Schluter (2021), Liang et al. (2022), Verma et al. (2023), Wang et al. (2021), Giannakopoulos et al. (2017), Ladhak et al. (2020a), Bhattacharjee et al. (2021), Tikhonov and Malykh (2022), Xiaorui (2023)
Swedish:
Varab and Schluter (2021), Aumiller et al. (2022), Tikhonov and Malykh (2022)
Tagalog:
Giannakopoulos et al. (2017)
Tajik:
Varab and Schluter (2021)
Tamil:
Hasan et al. (2021), Urlana et al. (2023), Varab and Schluter (2021), Liang et al. (2022), Verma et al. (2023), Bhattacharjee et al. (2021), Xiaorui (2023)
Telugu:
Hasan et al. (2021), Urlana et al. (2023), Urlana et al. (2022), Varab and Schluter (2021), Liang et al. (2022), Verma et al. (2023), Bhattacharjee et al. (2021), Xiaorui (2023)
Tetum:
Varab and Schluter (2021)
Thai:
Hasan et al. (2021), Palen-Michel and Lignos , Varab and Schluter (2021), Liang et al. (2022), Giannakopoulos et al. (2017), Ladhak et al. (2020a), Bhattacharjee et al. (2021), Xiaorui (2023)
Tibetan:
Palen-Michel and Lignos , Varab and Schluter (2021)
Tigrinya:
Hasan et al. (2021), Palen-Michel and Lignos , Varab and Schluter (2021), Liang et al. (2022), Bhattacharjee et al. (2021), Xiaorui (2023)
Turkish:
Hasan et al. (2021), Palen-Michel and Lignos , Scialom et al. (2020), Varab and Schluter (2021), Liang et al. (2022), Wang et al. (2021), Giannakopoulos et al. (2017), Ladhak et al. (2020a), Bhattacharjee et al. (2021), Xiaorui (2023)
Ukrainian:
Hasan et al. (2021), Palen-Michel and Lignos , Galeshchuk (2023), Varab and Schluter (2021), Liang et al. (2022), Verma et al. (2023), Wang et al. (2021), Giannakopoulos et al. (2017), Bhattacharjee et al. (2021), Tikhonov and Malykh (2022), Xiaorui (2023)
Urdu:
Hasan et al. (2021), Urlana et al. (2023), Palen-Michel and Lignos , Faheem et al. (2024), Varab and Schluter (2021), Liang et al. (2022), Verma et al. (2023), Bhattacharjee et al. (2021), Xiaorui (2023)
Uzbek:
Hasan et al. (2021), Palen-Michel and Lignos , Varab and Schluter (2021), Liang et al. (2022), Bhattacharjee et al. (2021), Xiaorui (2023)
Vietnamese:
Hasan et al. (2021), Palen-Michel and Lignos , Nguyen et al. (2019), Tran et al. (2020), Varab and Schluter (2021), Liang et al. (2022), Wang et al. (2021), Ladhak et al. (2020a), Bhattacharjee et al. (2021), Tikhonov and Malykh (2022), Xiaorui (2023)
Welsh:
Hasan et al. (2021), Varab and Schluter (2021), Liang et al. (2022), Bhattacharjee et al. (2021), Xiaorui (2023)
West African Pidgin English:
Hasan et al. (2021), Liang et al. (2022), Bhattacharjee et al. (2021), Xiaorui (2023)
Xhosa:
Varab and Schluter (2021)
Yoruba:
Hasan et al. (2021), Varab and Schluter (2021), Liang et al. (2022), Bhattacharjee et al. (2021), Xiaorui (2023)
Swahili:
Hasan et al. (2021), Palen-Michel and Lignos , Nguyen and Daumé III (2019), Varab and Schluter (2021), Liang et al. (2022), Bhattacharjee et al. (2021), Xiaorui (2023)